OpenStudy (anonymous):

PLEASE HELP!!!! MEDAL WILL BE REWARDED!!! Find the rectangular coordinates of the point with the polar coordinates (-7, 5pi/3) I keep getting (-6.97081,-0.63882) but i need the answer in fraction form...

OpenStudy (harsimran_hs4):

decimal to fraction -6.97081 = 697081/100000 now if there is any common factor of numerator and denominator in the fraction divide with it to get the simplified fraction that`s it

OpenStudy (anonymous):

It gives me such a high fraction though and i know thats not the answer! I want to say it (7/2, -7/2) but im not sure....

OpenStudy (harsimran_hs4):

no the answer should be x = -7 * cos(300 degree) or -7 * cos(5 pi/3) y = -7 * sin(300 degree) or -7 * sin(5 pi/3)

OpenStudy (anonymous):

\[(-7\cos(\frac{5\pi}{3}), -7\sin(\frac{5\pi}{3}))\]

OpenStudy (anonymous):

Those arnt the answers, i only havve 4 choices, here they are (-7/2, 7/2) (7/2, -7/2) (7(radical3)/2, -7/2) (-7/2, 7(radical3)/2)

OpenStudy (anonymous):

i know they aren't the answers that is the way you find the answers however your job is to find \(\cos(\frac{5\pi}{3})\) and \(\sin(\frac{5\pi}{3})\) and then multiply the result by \(-7\)

OpenStudy (harsimran_hs4):

well we want that you put the values of sin300 and cos300 to get the answers

OpenStudy (anonymous):

forget degrees, this problem has nothing to do with degrees

OpenStudy (anonymous):

ohh okay, hold on...i get the third answer on that list...is that correct? or am i still messing up!?

OpenStudy (harsimran_hs4):

i think 4th should be correct tell us the values you found out for sin(5pi/3) and cos(5pi/3)

OpenStudy (anonymous):

\(\cos(\frac{5\pi}{3})=\frac{1}{2}\)

OpenStudy (anonymous):

Ohhh okay, i see how to work the problem now. I wasnt typing in my values right. I was doing them opposite from what it was asking, so yes it is the 4th answer im pretty sure